# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v1

@@ II.2 Partial Rankings
II.3 Optimization
We propose a fast nonparametric agglomerative algorithm based on a alternating optimisation strategy to approximate the MAP estimates $\bm{\hat{r}}$, $\bm{\hat{\sigma}}$ for the node rankings and strengths. Taking the gradient to optimize $\mathcal{L}$ with respect to $\bm{\sigma}$, while keeping the rank vector $\bm{r}$ fixed, yields
@@ II.3 Optimization
Eq. 19 defines a set of $R$ self-consistent equations for the $\sigma_{r}$’s which can be iteratively solved until convergence. If $E(\bm{r})$ is the number of non-zero entries in the matrix $\bm{\omega}$ when the ranks are equal to $\bm{r}$, then the complexity of this algorithm is $O(E(\bm{r}))$. These iterative updates for the player’s strengths have been previously studied in newman2023efficient , where it was shown that they are guaranteed to converge and do so quite efficiently in practice.
To update $\bm{r}$ given $\bm{\sigma}$, we can greedily identify the pair of consecutive ranks $r<r^{\prime}$ that produces the largest decrease $\Delta\mathcal{L}(r,r^{\prime})$ in the negative log-posterior when the two ranks are merged. Let $R$ represent the current number of ranks, and let $\bm{n}$ denote the sizes of the rank groups before the merge. The change in the negative log-posterior resulting from merging $r$ and $r^{\prime}$ is given by:
@@
where $(r,r^{\prime})$ represents the new rank group formed by merging groups $r,r^{\prime}$ and
$C(R)=\log N+\log\binom{N-1}{R-1}+\log N!,$ (21)
$f(r,r^{\prime})=\sum_{r,r^{\prime}}\omega_{r,r^{\prime}}\log\left[\frac{\sigma_{r}+\sigma_{r^{\prime}}}{\sigma_{r}}\right].$ (23)
$C(R)$ accounts for the contribution to the negative log-posterior given by the number of rankings $R$, $g(r)$ describes contributions from group sizes and strengths, and $f(r,r^{\prime})$ captures the interaction terms between the rank groups based on the comparison matrix $\bm{\omega}$ among the rank groups.
@@
Having computed the strength for the merged cluster $(r,r^{\prime})$, we can repeat the process and do another full update for the strengths $\bm{\sigma}$ given the new ranking $\bm{r}$ after the merge. Again, this requires using the previous iterative update with complexity $O(E(\bm{r}))$. We can then continue alternating the updates $\bm{\sigma}|\bm{r}$ and $\bm{r}|\bm{\sigma}$ until all groups have been merged, at which point we can inspect all the examined rankings and select the ranking $\bm{\hat{r}}$ associated with the highest posterior probability (the lowest value of $\mathcal{L}(\bm{r},\bm{\sigma})$).
@@
where $N$ is the number of nodes or players in the network, $\bm{r}$ is the optimal rank vector for each value $R$ of the number of unique ranks during the merge process, and $\alpha\in[0,1]$ depends on the density of the matrix $\bm{\omega}$ as the ranks are progressively merged. In the worst-case scenario, where the matrix $\bm{\omega}$ is maximally dense ($E(\bm{r})=R^{2}$ at every step), the complexity would scale as $O(N^{3})$. However, such maximal density is highly unlikely for $R\approx N$, especially in real, sparse networks. In practice, we observe a runtime scaling of approximately $O(N^{2})$ with the network size $N$; see Fig. 6 in Appendix A. We note that this algorithm is greedy in nature, and therefore is only guaranteed to identify a local optimum for the MAP estimates $\bm{\hat{r}}$, $\bm{\hat{\sigma}}$. However, exact optimization over node groupings is likely NP-hard (as is the case with many other clustering problems welch1982algorithmic ), making exact inference intractable in all but the smallest of networks. Greedy agglomerative algorithms have been shown to closely approximate the optimal log posterior probabilities obtained through exact enumeration and simulated annealing for other network clustering tasks in which we expect to preserve spatial contiguity kirkley2022spatial ; morel2024bayesian or an initial temporal ordering kirkley2024inference , similar in nature to preserving the initial ordering of the BT scores under the final inferred partial rankings, which is overwhelmingly the case in practice.
Fig. 1 shows the results of applying both the Bradley-Terry (BT) model and our partial rankings model (with the Bradley Terry likelihood of Eq. 3) to a small dataset of dominance interactions in a pack of wolves van198711 . We notice that our model offers a more concise interpretation of the data by grouping individuals with similar BT scores into the same rankings.
@@ III.1 Synthetic Match Datasets
Figure 2: Partial rankings in synthetic datasets. (a) Heatmap of the number of rankings $R$ inferred by the partial rankings model, as a function of the parameters $(\sigma,\langle k\rangle)$ of the synthetic model in Sec. III.1. (b) Heatmap of the difference in the total number of times the partial rankings and Bradley-Terry models emerged as the best-performing model (in terms of the posterior odds ratio Eq. 27), out of $20$ trials of synthetically generated networks. Positive values indicate that the partial rankings model more often emerged as the best performer, and negative values indicate that the Bradley-Terry model was more often preferred. (c) Heatmap of the log posterior odds ratio (Eq. 27) between the Bradley-Terry and the partial rankings model across the simulations. Positive values indicate a preference for the partial rankings model, and negative values a preference for the Bradley-Terry model. (d) The top panel displays the average number of rankings inferred by the best-performing model (blue) and our partial rankings model (red) as a function of the average degree $\langle k\rangle$ of the network of matches for the case in which no planted partial rankings are present. The bottom panel displays the log posterior odds ratio (normalized per node) between the two models. Negative values of this difference indicate a preference for the Bradley-Terry model and positive values a preference for the partial rankings model. All results were obtained by averaging over $20$ different simulations from the synthetic network model of Sec. III.1, and error bars indicate $2$ standard errors in the mean.
@@
    For a chosen number of matches $M$, randomly select pairs of players (with repetition) and simulate match outcomes using the Bradley-Terry model.
@@
By varying $M$ and $\sigma$, we generate networks of matches with different levels of sparsity and strength separation, allowing us to assess the performance of our algorithm across a wide range of scenarios. Results, shown in Fig. 2, depict the algorithm’s ability to recover the planted rankings as a function of $\sigma$ and the average degree of the network $\langle k\rangle=M/N$. For each $(\sigma,\langle k\rangle)$ pair, $20$ different networks were generated, and the results were then averaged.
@@
As expected, we observe that for low values of $\langle k\rangle$, there is not enough evidence to recover the planted rankings regardless of the strength separation between players. In such sparse conditions, the algorithm assigns all players to the same rank, as shown in Fig. 2(a). This behaviour contrasts sharply with the Bradley-Terry (BT) model, which still produces a complete ranking of the nodes even in these highly sparse conditions. As the number of matches increases and the networks become denser, the model gains sufficient statistical evidence to distinguish between the rankings and eventually recovers the planted structure. Notably, the number of matches required to correctly recover the rankings decreases rapidly with greater strength separation between players. In these cases, stronger players consistently outperform weaker ones, creating clearer signals that enable the ranking structure to be resolved with less data.
@@
Perhaps more notable is the fact that our model does not consistently outperform the Bradley-Terry model in scenarios where players are nearly equal in strength. When the separation in player strengths is small, our partial rankings model outperforms BT in sparse networks, as shown in Figs. 2(b)-(c). In these cases, the PR model effectively leverages its ability to group players into shared ranks, avoiding overfitting. However, as the networks become denser, while our model accurately recovers the planted rankings even in this low-separation regime, it does not offer a more parsimonious description of the data (in terms of the posterior odds ratio of Eq. 27) compared to the standard Bradley-Terry model which, in contrast, infers a fully ranked structure for the nodes with $R=N=50$. This suggests a noisy regime where the evidence supporting partial rankings is relatively weak. The Bradley-Terry likelihood provides a more flexible fit for fine-gained variations in the data in these dense regimes (effectively overfitting the data), even if its prior poorly reflects the underlying data.
In Fig. 2(d), we examine the behaviour of both models in a scenario where no partial rankings are present, and the networks consist of $N=150$ players, each with its own unique score. The results, averaged over 20 network realisations, show that for small values of $\langle k\rangle$, our partial rankings model still provides a more parsimonious description of the data. However, as $\langle k\rangle$ increases, we observe a sharp transition where the BT model rapidly provides a better description of the data. We also observe that, as $\langle k\rangle$ grows, the PR model gradually infers more rankings, albeit extremely slowly, suggesting that, in principle, the PR model can recover the full ranking given enough data. Interestingly, as the inferred number of groups increases with $\langle k\rangle$, the log posterior odds ratio, defined as
$\Delta\mathcal{L}=\log\frac{P_{BT}(\bm{\pi|A})}{P_{PR}(\bm{\pi|A})},$ (27)
@@
decreases. This is somewhat counterintuitive, as one might expect the posterior odds ratio to improve as the number of rankings inferred by the PR model approaches the true number of rankings. However, as mentioned in Sec. II, the PR model tends to heavily penalise the presence of a large number of rankings so that the net effect of increasing the number of rankings is a reduction in the posterior odds ratio, with the PR model becoming a less efficient encoding for the rankings as the number of inferred groups grows. In practice, the best practice is to evaluate both an original ranking method and our corresponding partial ranking method to determine which provides a more parsimonious fit to the data in terms of log-posterior odds.
@@
To assess the goodness of the recovered partitions, we compute the Kendall rank correlation coefficient ($\tau_{B}$), adjusted for ties, between the ground truth rankings and the rankings inferred by the BT and PR models. This comparison is performed both when partial rankings are present and when they are not. We observe that when partial rankings are present in the network, the PR model can quickly attain perfect recovery as the networks become denser. On the other hand, while the BT model still achieves high $\tau_{B}$ scores, it never achieves perfect ordinal association and always underperforms relative to the PR model. However, when no partial rankings are present, and each player is assigned a unique score, the roles are reversed, and BT always displays a higher ordinal association with the ground truth rankings (see Appendix B).
@@
We also compare our model with the partial rankings derived by applying 1D mean shift (MS) clustering comaniciu2002mean to the rankings inferred by the BT model. We observe that our algorithm can always recover the correct number of partial rankings when these are present, while MS fails to do so. More importantly, our algorithm can adjust the number of inferred clusters as more data becomes available, even in cases when no partial rankings exist, resulting in a more accurate description of the rankings. In contrast, mean shift clustering consistently infers roughly the same number of partial rankings regardless of the evidence provided by the data (see Appendix C).
@@ III.2 Real Match Datasets
Figure 3: Partial rankings in real networks of pairwise comparisons. (a) (top) Rescaled number of ranks per node inferred by the best-performing model and (bottom) log posterior odds ratio per node, both as a function of the average degree for all real-world networks considered in the study (see Table 1). The colours of the points indicate the different categories to which the datasets belong, while the shape of the markers indicates which model (PR or BT) emerged as the best-performing method (in terms of the posterior odds ratio). Positive values of the log posterior odds ratio indicate evidence in favour of the PR model; negative values indicate evidence in favour of the BT model. (b) Effective number of ranks, $R^{*}$ (Eq. 28), as a function of the number of unique ranks $R$ inferred by the PR algorithm. The black dashed line represents the line $R^{*}=R$. (c) Number of rankings inferred via the partial rankings algorithm as a function of the number of rankings inferred via mean shift clustering. Point colours indicate the $\tau_{B}$ score between the two inferred rankings. The black dashed line represents the line $R_{PR}=R_{MS}$. (d) Number of nodes of each network as a function of the number of rankings $R$ inferred by the partial rankings algorithm. Point colours indicate the $\tau_{B}$ score between the rankings inferred via the partial rankings algorithm and those inferred by the Bradley-Terry model. The dashed black line indicates the line $R=N$.
@@
Table 1: Datasets analysed using the partial rankings algorithm and Bradley-Terry algorithm, in order of decreasing log posterior odds ratio in Eq. 27 (i.e. increasingly in favor of the original Bradley-Terry model). $N$, $M$, and $\langle k\rangle$ indicate the number of nodes, edges, and average degree of each network, respectively. $R$ and $R^{*}$ indicate the number of partial rankings inferred via the PR algorithm and the effective number of partial rankings (Eq. 28). $\mathcal{L}$ indicates the logarithm of the posterior odds ratio between the PR and BT models (Eq. 27).
In this section we focus on how our partial rankings algorithm may be used to derive meaningful patterns in real-world data. We consider the datasets compiled in jerdee2024luck and available at jerdeegit . In addition, we also consider a dataset of social hierarchies among a pack of wolves van198711 ; newman2023efficient , and other sets of directed networks available at netzschleuder . The datasets are summarized in Table 1.
We apply our partial rankings algorithm and the Bradley-Terry model to all match lists in the dataset and compare the results. In Fig. 3(a), we show the number of inferred ranks per node and the value of the log posterior odds ratio per node as a function of the average degree of each network. To standardise the number of ranks per node within the interval $[0,1]$, we rescale it as $(R-1)/(N-1)$, where $N$ is the number of nodes in the network and $R$ is the total number of ranks inferred by the best-performing model. This quantity will be $1$ for perfect orderings in which $R=N$. It will be 00 if no rankings are inferred and all nodes are assigned the same strength, and it will take intermediate values when non-trivial partial rankings are present. The data reveals three distinct phases based on the density of the underlying network. For low average degree values, $\langle k\rangle$, the evidence is insufficient to support any meaningful ranking, leading to all actors being assigned the same strength. As connectivity increases, partial rankings begin to offer a more parsimonious description of the data compared to a perfect ordering of the nodes. In this intermediate regime, non-trivial subdivisions emerge, where actors are grouped by strength: all nodes within a group share the same strength, but strengths vary across groups. Finally, as the networks become sufficiently dense, enough data is available to support a perfect ordering of the nodes, and the Bradley-Terry model provides a more parsimonious description of the data.
@@
where $n_{r}$ is the size of rank $r$ and $N$ is the total number of nodes in the network. This measure reflects the balance in the size distribution of rankings: $R^{*}$ will achieve a maximum value of $R$ when all the rankings are of equal size. In contrast, it approaches a minimal value close to one when the distribution is highly imbalanced, with most nodes concentrated in a single rank. Fig. 3(b) shows the $R^{*}$ as a function of the inferred partial rankings $R$ for the PR model. While for most datasets $R^{*}\simeq R$, indicating that in most cases the PR algorithm infers roughly equally sized rankings, there are some notable exceptions, most notably the dataset of ATP tennis matches between 2010 and 2019, which displays a considerably smaller effective number of groups, reflecting a pronounced imbalance in the sizes of the inferred ranks. Specifically, the strongest 15 players are distributed across three partial rankings, while over $84\%$ of all players are grouped into the weakest rank. This skewed decomposition likely stems from the structure of professional tennis tournaments. ATP matches are generally organised in a knock-out format, where losers are progressively eliminated. While this structure is efficient and straightforward, it can result in noisy rankings, as strong players may be eliminated early due to random fluctuations or unfavourable matchups. To mitigate this, the ATP organises multiple knock-out tournaments throughout the year, ensuring that the strongest players consistently rise to the top despite potential variability in individual tournaments. However, this system disproportionately concentrates matches among the strongest players, who predominantly compete against one another as they progress through the brackets. Consequently, the dataset contains ample information to distinguish the strongest players but far less data to differentiate among weaker players, leading to highly unbalanced partial rankings in terms of size.
@@
Again, we compare the results obtained via our PR algorithm with those obtained by applying 1D Mean Shift clustering to the rankings inferred by the Bradley-Terry model. In Fig. 3(c), we plot the number of partial rankings, $R_{PR}$, inferred by the PR algorithm against the number of partial rankings, $R_{MS}$, inferred by Mean Shift. While a general positive correlation is observed between the two quantities, there is significant variability in the number of rankings inferred by the two methods. Moreover, even when the total number of rankings inferred by both algorithms is similar, the rankings themselves may differ substantially, as evidenced by the Kendall rank correlation ($\tau_{B}$) scores also shown in Fig. 3(c). Specifically, $\tau_{B}$ scores appear to increase with the total number of groups inferred by both algorithms, regardless of how close $R_{PR}$ and $R_{MS}$ are. This suggests that the increase in $\tau_{B}$ scores is likely an artefact resulting from the fact that the algorithms are likely to align on most pairwise relationships as the number of groups grows rather than reflecting a similarity in the group structures inferred by the two approaches.

### v2

@@ II.3 Bayesian Nonparametric Model
II.4 Optimization
{algorithm}
Greedy Algorithm for MAP inference of Partial Rankings
1:Initialize ranking vector $r$ with each node in its own group (i.e. $R=N$)
2:Initialize strength vector $\sigma$ with initial BT estimates
3:$L_{\text{current}}\leftarrow\text{ComputeLoss}(r,\sigma)$
4:$L_{\text{best}}\leftarrow L_{\text{current}}$
5:$r_{\text{best}}\leftarrow r$, $\sigma_{\text{best}}\leftarrow\sigma$
6:while $R>1$ do $\triangleright$ Continue merging until a single group remains
7:     /* Evaluate potential merges between adjacent groups */
8:     for each adjacent pair $(i,i+1)$ in $r$ do
9:         $\Delta L[i]\leftarrow\text{ComputeDeltaL}(\text{MergeGroup}(i,i+1),r,\sigma)$
10:     end for
11:     $(i_{\text{merge}},j_{\text{merge}})\leftarrow\arg\min_{(i,i+1)}\Delta L[i]$
12:     /* Merge the selected groups */
13:     $r\leftarrow\text{MergeGroups}(r,i_{\text{merge}},j_{\text{merge}})$
14:     $\sigma_{\text{merged}}\leftarrow\text{SolveForMergedSigma}(i_{\text{merge}},j_{\text{merge}},r,\sigma)$
15:     Update $\sigma$ to incorporate the merged group and adjust the remaining strengths accordingly
16:     $L_{\text{current}}\leftarrow\text{ComputeLoss}(r,\sigma)$
17:     if $L_{\text{current}}<L_{\text{best}}$ then
18:         $L_{\text{best}}\leftarrow L_{\text{current}}$
19:         $r_{\text{best}}\leftarrow r$, $\sigma_{\text{best}}\leftarrow\sigma$
20:     end if
21:     $R\leftarrow$ number of groups in $r$
22:end while
23:return $r_{\text{best}},\sigma_{\text{best}}$
We propose a fast nonparametric agglomerative algorithm based on an alternating optimization strategy to approximate the MAP estimates $\bm{\hat{r}}$, $\bm{\hat{\sigma}}$ for the node rankings and strengths. Starting from an initial condition in which each node belongs to its own group, so that $R=N$, the algorithm alternates between updating the unique strengths $\bm{\sigma}$ and rankings $\bm{r}$ so as to minimize the negative log-posterior $\mathcal{L}$ (Eq. II.3).
Taking the gradient to optimize $\mathcal{L}$ with respect to $\bm{\sigma}$, while keeping the rank vector $\bm{r}$ fixed, yields
@@ II.4 Optimization
Eq. 28 defines a set of $R$ self-consistent equations for the $\sigma_{r}$’s which can be iteratively solved until convergence. If $E(\bm{r})$ is the number of non-zero entries in the matrix $\bm{\omega}$ when the ranks are equal to $\bm{r}$, then the complexity of this algorithm is $O(E(\bm{r}))$. These iterative updates for the player’s strengths are of the same form as those introduced in Eq. 12 and benefit from the same convergence guarantees.
To update $\bm{r}$ given $\bm{\sigma}$, we can greedily identify the pair of consecutive ranks $r<r^{\prime}$ that produces the largest decrease $\Delta\mathcal{L}(r,r^{\prime})$ in the negative log-posterior when the two ranks are merged. To do this, it is useful to rewrite the negative log-posterior as follows:
$\mathcal{L}(\bm{r})=C(R)+\sum_{r=1}^{R}g(r)+\sum_{r,r^{\prime}=1}^{R}f(r,r^{\prime})$ (29)
where
$C(R)=\log N+\log\binom{N-1}{R-1}+\log N!,$ (30)
$f(r,r^{\prime})=\sum_{r,r^{\prime}}\omega_{r,r^{\prime}}\log\left[\frac{\sigma_{r}+\sigma_{r^{\prime}}}{\sigma_{r}}\right]$ (32)
and we have dropped the dependence on the strength parameters as we consider them fixed. $C(R)$ accounts for the contribution to the negative log-posterior given by the number of rankings $R$, $g(r)$ describes contributions from group sizes and strengths, and $f(r,r^{\prime})$ captures the interaction terms between the rank groups based on the comparison matrix $\bm{\omega}$ among the rank groups.
Let $R$ represent the current number of ranks, and let $\bm{n}$ denote the sizes of the rank groups before the merge. Then the change in the negative log-posterior resulting from merging $r$ and $r^{\prime}$ is given by:
@@
where $(r,r^{\prime})$ represents the new rank group formed by merging groups $r,r^{\prime}$.
@@
Having computed the strength for the merged cluster $(r,r^{\prime})$, we can repeat the process and do another full update for the strengths $\bm{\sigma}$ given the new ranking $\bm{r}$ after the merge. Again, this requires using the previous iterative update with complexity $O(E(\bm{r}))$. We can then continue alternating the updates $\bm{\sigma}|\bm{r}$ and $\bm{r}|\bm{\sigma}$ until all groups have been merged, at which point we can inspect all the examined rankings and select the ranking $\bm{\hat{r}}$ associated with the highest posterior probability (the lowest value of $\mathcal{L}(\bm{r},\bm{\sigma})$). Pseudocode for the algorithm is shown in Algorithm II.4.
Figure 1: Partial rankings in a small example network. (a) Rankings inferred by the BT model and (b) the PR algorithm, for a dataset capturing hierarchical relationships in a pack of wolves. The nodes are labeled according to their inferred ranking. The distance along the y-axis and the node colors are proportional to the inferred strength $\pi_{i}$ of each node $i$, with the strongest nodes placed at the top. In this case, there was not enough statistical evidence in the edges to justify separating the node ranks $\{2,3\}$, $\{6,7,8,9\}$, and $\{11,12\}$ on the left hand side, so the partial ranking method grouped these nodes together into the same partial rankings.
@@
where $N$ is the number of nodes or players in the network, $\bm{r}$ is the optimal rank vector for each value $R$ of the number of unique ranks during the merge process, and $\alpha\in[0,1]$ depends on the density of the matrix $\bm{\omega}$ as the ranks are progressively merged. In the worst-case scenario, where the matrix $\bm{\omega}$ is maximally dense ($E(\bm{r})=R^{2}$ at every step), the complexity would scale as $O(N^{3})$. However, such maximal density is highly unlikely for $R\approx N$, especially in real, sparse networks. In practice, we observe a runtime scaling of approximately $O(N^{2})$ with the network size $N$; see Fig. 6 in Appendix B. We note that this algorithm is greedy in nature, and therefore is only guaranteed to identify a local optimum for the MAP estimates $\bm{\hat{r}}$, $\bm{\hat{\sigma}}$. However, exact optimization over node groupings is likely NP-hard (as is the case with many other clustering problems welch1982algorithmic ), making exact inference intractable in all but the smallest of networks. Greedy agglomerative algorithms have been shown to closely approximate the optimal log posterior probabilities obtained through exact enumeration and simulated annealing for other network clustering tasks in which we expect to preserve spatial contiguity kirkley2022spatial ; morel2024bayesian or an initial temporal ordering kirkley2024inference , similar in nature to preserving the initial ordering of the BT scores under the final inferred partial rankings, which is overwhelmingly the case in practice.
Fig. 1 shows the results of applying both the BT model and our Partial Rankings (PR) algorithm (with the BT likelihood of Eq. 3) to a small dataset of dominance interactions in a pack of wolves van198711 . We notice that our model offers a more concise interpretation of the data by grouping individuals with similar BT scores into the same rankings.
@@ III.1 Synthetic Match Datasets
Figure 2: Partial rankings in synthetic datasets. (a) Heatmap of the number of rankings $R$ inferred by the partial rankings model, as a function of the parameters $(\sigma,\langle k\rangle)$ of the synthetic model in Sec. III.1. (b) Heatmap of the log posterior odds ratio (Eq. 37) between the BT and the partial rankings model across the simulations. Positive values indicate a preference for the partial rankings model, and negative values a preference for the BT model. (c) The top panel displays the average number of rankings inferred by the best-performing model (blue) and our partial rankings model (red) as a function of the average degree $\langle k\rangle$ of the network of matches for the case in which no planted partial rankings are present. The bottom panel displays the log posterior odds ratio (normalized per node) between the two models. Negative values of this difference indicate a preference for the BT model and positive values a preference for the partial rankings model. All results were obtained by averaging over $20$ different simulations from the synthetic network model of Sec. III.1, and error bars indicate $2$ standard errors in the mean. Except near the transition point, error bars are smaller than the marker size.
@@
    For a chosen number of matches $M$, randomly select pairs of players (with repetition) and simulate match outcomes using the BT model.
@@
By varying $M$ and $\sigma$, we generate networks of matches with different levels of sparsity and strength separation, allowing us to assess the performance of our algorithm across a wide range of scenarios. Results, shown in Fig. 2, depict the algorithm’s ability to recover the planted rankings as a function of $\sigma$ and the average degree of the network $\langle k\rangle=M/N$. For each $(\sigma,\langle k\rangle)$ pair, $20$ different networks were generated, and the results were then averaged. The ranges of values for $\langle k\rangle$ and $\sigma$ were chosen so as to capture all the relevant behaviors of the model.
@@
As expected, we observe that for low values of $\langle k\rangle$, there is not enough evidence to recover the planted rankings regardless of the strength separation between players. In such sparse conditions, the algorithm assigns all players to the same rank, as shown in Fig. 2(a). This behavior contrasts sharply with the BT model, which still produces a complete ranking of the nodes even in these highly sparse conditions. As the number of matches increases and the networks become denser, the model gains sufficient statistical evidence to distinguish between the rankings and eventually recovers the planted structure. Notably, the number of matches required to correctly recover the rankings decreases rapidly with greater strength separation between players. In these cases, stronger players consistently outperform weaker ones, creating clearer signals that enable the ranking structure to be resolved with less data.
Perhaps more notable is the fact that our model does not consistently outperform the BT model in scenarios where players are nearly equal in strength. Fig. 2(b) shows a heatmap of the log posterior odds ratio, defined as
$\Delta\mathcal{L}=\log\frac{P_{PR}(\bm{\pi|A})}{P_{BT}(\bm{\pi|A})},$ (37)
@@
as a function of $\langle k\rangle$ and $\sigma$. We observe that, when the separation in player strengths is small, our partial rankings model outperforms BT in sparse networks. In these cases, the PR model effectively leverages its ability to group players into shared ranks, avoiding overfitting. However, as the networks become denser, while our model accurately recovers the planted rankings even in this low-separation regime, it does not offer a more parsimonious description of the data (in terms of the posterior odds ratio of Eq. 37) compared to the standard BT model which, in contrast, infers a fully ranked structure for the nodes with $R=N=50$. This suggests a noisy regime where the evidence supporting partial rankings is relatively weak. The BT likelihood provides a more flexible fit for fine-gained variations in the data in these dense regimes (effectively overfitting the data), even if its prior poorly reflects the underlying data.
In Fig. 2(c), we analyze the performance of both models in the absence of partial rankings, using networks of $N=150$ players, each assigned a unique score sampled from a logistic distribution with mean zero and scale one—a setting in which the BT model is expected to accurately capture the data. Averaged over 20 network realizations, the results show that for small values of $\langle k\rangle$, our partial rankings (PR) model offers a more parsimonious fit, as the limited number of interactions prevents reliable inference of a complete ranking. As $\langle k\rangle$ increases, however, we observe a sharp transition in which the BT model quickly surpasses the PR model in explanatory power. We also observe that the PR model continues to infer more rankings as $\langle k\rangle$ grows, albeit very slowly—indicating that, in principle, it is capable of recovering the full ranking given sufficient data.
@@
Interestingly, as the inferred number of groups increases with $\langle k\rangle$, the log posterior odds ratio decreases. This is somewhat counterintuitive, as one might expect the posterior odds ratio to improve as the number of rankings inferred by the PR model approaches the true number of rankings. However, as mentioned in Sec. II, the PR model tends to heavily penalize the presence of a large number of rankings so that the net effect of increasing the number of rankings is a reduction in the posterior odds ratio, with the PR model becoming a less efficient encoding for the rankings as the number of inferred groups grows. In practice, the best practice is to evaluate both an original ranking method and our corresponding partial ranking method to determine which provides a more parsimonious fit to the data in terms of log-posterior odds.
@@
To assess the goodness of the recovered partitions, in Appendix C we compute the Kendall rank correlation coefficient ($\tau_{B}$), adjusted for ties, between the ground truth rankings and the rankings inferred by the BT and PR models. This comparison is performed both when partial rankings are present and when they are not. We observe that when partial rankings are present in the network, the PR model can quickly attain perfect recovery as the networks become denser. On the other hand, while the BT model still achieves high $\tau_{B}$ scores, it never achieves perfect ordinal association and always underperforms relative to the PR model. However, when no partial rankings are present, and each player is assigned a unique score, the roles are reversed, and BT always displays a higher ordinal association with the ground truth rankings.
@@
We also compare our model with the partial rankings derived by applying 1D mean shift (MS) clustering comaniciu2002mean to the rankings inferred by the BT model, in Appendix D. We observe that our algorithm can always recover the correct number of partial rankings when these are present, while MS fails to do so. More importantly, our algorithm can adjust the number of inferred clusters as more data becomes available, even in cases when no partial rankings exist, resulting in a more accurate description of the rankings. In contrast, mean shift clustering consistently infers roughly the same number of partial rankings regardless of the evidence provided by the data.
@@ III.2 Real Match Datasets
Figure 3: Partial rankings in real networks of pairwise comparisons. (a) (top) Rescaled number of ranks per node inferred by the best-performing model and (bottom) log posterior odds ratio per node, both as a function of the average degree for all real-world networks considered in the study (see Table 1). The colours of the points indicate the different categories to which the datasets belong, while the shape of the markers indicates which model (PR or BT) emerged as the best-performing method (in terms of the posterior odds ratio). Positive values of the log posterior odds ratio indicate evidence in favour of the PR model; negative values indicate evidence in favour of the BT model. (b) Effective number of ranks, $R^{*}$ (Eq. 38), as a function of the number of unique ranks $R$ inferred by the PR algorithm. The black dashed line represents the line $R^{*}=R$. (c) Number of rankings inferred via the partial rankings algorithm as a function of the number of rankings inferred via mean shift clustering. Point colours indicate the $\tau_{B}$ score between the two inferred rankings. The black dashed line represents the line $R_{PR}=R_{MS}$. (d) Number of nodes of each network as a function of the number of rankings $R$ inferred by the partial rankings algorithm. Point colours indicate the $\tau_{B}$ score between the rankings inferred via the PR algorithm and those inferred by the BT model. The dashed black line indicates the line $R=N$.
@@
Table 1: Datasets analysed using the PR and BT algorithms, in order of decreasing log posterior odds ratio in Eq. 37 (i.e. increasingly in favor of the original BT model). $N$, $M$, and $\langle k\rangle$ indicate the number of nodes, edges, and average degree of each network, respectively. $R$ and $R^{*}$ indicate the number of partial rankings inferred via the PR algorithm and the effective number of partial rankings (Eq. 38). $\mathcal{L}$ indicates the logarithm of the posterior odds ratio between the PR and BT models (Eq. 37).
In this section, we focus on how our PR algorithm may be used to derive meaningful patterns in real-world data. We consider the datasets compiled in jerdee2024luck and available at jerdeegit . In addition, we also consider a dataset of social hierarchies among a pack of wolves van198711 ; newman2023efficient , and other sets of directed networks available at netzschleuder . The datasets are summarized in Table 1.
We apply our PR algorithm and the BT model to all match lists in the dataset and compare the results. In Fig. 3(a), we show the number of inferred ranks per node and the value of the log posterior odds ratio per node as a function of the average degree of each network. To standardize the number of ranks per node within the interval $[0,1]$, we rescale it as $(R-1)/(N-1)$, where $N$ is the number of nodes in the network and $R$ is the total number of ranks inferred by the best-performing model. This quantity will be $1$ for perfect orderings in which $R=N$. It will be 00 if no rankings are inferred and all nodes are assigned the same strength, and it will take intermediate values when non-trivial partial rankings are present. The data reveals three distinct phases based on the density of the underlying network. For low average degree values, $\langle k\rangle$, the evidence is insufficient to support any meaningful ranking, leading to all actors being assigned the same strength. As connectivity increases, partial rankings begin to offer a more parsimonious description of the data compared to a perfect ordering of the nodes. In this intermediate regime, non-trivial subdivisions emerge, where actors are grouped by strength: all nodes within a group share the same strength, but strengths vary across groups. Finally, as the networks become sufficiently dense, enough data is available to support a perfect ordering of the nodes, and the BT model provides a more parsimonious description of the data.
@@
where $n_{r}$ is the size of rank $r$ and $N$ is the total number of nodes in the network. This measure reflects the balance in the size distribution of rankings: $R^{*}$ will achieve a maximum value of $R$ when all the rankings are of equal size. In contrast, it approaches a minimal value close to one when the distribution is highly imbalanced, with most nodes concentrated in a single rank. Fig. 3(b) shows the $R^{*}$ as a function of the inferred partial rankings $R$ for the PR model. While for most datasets $R^{*}\simeq R$, indicating that in most cases the PR algorithm infers roughly equally sized rankings, there are some notable exceptions, most notably the dataset of ATP tennis matches between 2010 and 2019, which displays a considerably smaller effective number of groups, reflecting a pronounced imbalance in the sizes of the inferred ranks. Specifically, the strongest 15 players are distributed across three partial rankings, while over $84\%$ of all players are grouped into the weakest rank. This skewed decomposition likely stems from the structure of professional tennis tournaments. ATP matches are generally organized in a knock-out format, where losers are progressively eliminated. While this structure is efficient and straightforward, it can result in noisy rankings, as strong players may be eliminated early due to random fluctuations or unfavorable matchups. To mitigate this, the ATP organizes multiple knock-out tournaments throughout the year, ensuring that the strongest players consistently rise to the top despite potential variability in individual tournaments. However, this system disproportionately concentrates matches among the strongest players, who predominantly compete against one another as they progress through the brackets. Consequently, the dataset contains ample information to distinguish the strongest players but far less data to differentiate among weaker players, leading to highly unbalanced partial rankings in terms of size.
@@
Again, we compare the results obtained via our PR algorithm with those obtained by applying 1D Mean Shift clustering to the rankings inferred by the BT model. In Fig. 3(c), we plot the number of partial rankings, $R_{PR}$, inferred by the PR algorithm against the number of partial rankings, $R_{MS}$, inferred by Mean Shift. While a general positive correlation is observed between the two quantities, there is significant variability in the number of rankings inferred by the two methods. Moreover, even when the total number of rankings inferred by both algorithms is similar, the rankings themselves may differ substantially, as evidenced by the Kendall rank correlation ($\tau_{B}$) scores also shown in Fig. 3(c). Specifically, $\tau_{B}$ scores appear to increase with the total number of groups inferred by both algorithms, regardless of how close $R_{PR}$ and $R_{MS}$ are. This suggests that the increase in $\tau_{B}$ scores is likely an artefact resulting from the fact that the algorithms are likely to align on most pairwise relationships as the number of groups grows rather than reflecting a similarity in the group structures inferred by the two approaches.